A high-yield bond has the following features:

__________________________________________________

Principal amount: $1,000

Interest rate (the coupon): 11.50%

Maturity: 10 years

Sinking fund: None

Call feature: After two years

Call penalty: One year’s interest

___________________________________________________

a.) If comparable yields are 12 percent, what should be the price of this bond?

b.) Would you expect the firm to call the bond if yields are 12 percent?

c.) If comparable yields are 8 percent, what should be the price of the bond?

d.) Would you expect the firm to call the bond today if yields are 8 percent?

e.) If you expected the bond to be called after three years, what is the maximum price you would pay for the bond if the current interest rate is 8 percent?
Answer #1

1.
Price=PV(12%,10,-11.5%*1000,-1000)
=$971.75

2.
Yield to call=RATE(2,11.5%*1000,-971.75,1000+11.5%*1000)=18.58%

The bond will not be called as yield to call is more than yield to maturity

3.
Price=PV(8%,10,-11.5%*1000,-1000)=$1,234.85

4.
Yield to call=RATE(2,11.5%*1000,-1234.85,1000+11.5%*1000)=4.57%

The bond will be called as yield to call is less than yield to maturity

5.

=PV(8%,3,-11.5%*1000,-1000-11.5%*1000)=$1,181.49